Which causes a greater effect to the magnitude of the velocity of an object in uniform circular motion: variation in centripetal force or variation in radius?


1
Expert's answer
2021-04-06T13:53:45-0400

The centripetal force can be found as follows:


"F_c=ma_c=m\dfrac{v^2}{r}."

From this formula we can find the velocity:


"v=\sqrt{\dfrac{F_cr}{m}}."

As we can see from the formula, the variation in centripetal force and variation in radius causes the same effect on the velocity.